Please write down the detailed steps3. Inimitably instinctive investigator Sheerluck Hopeless tests her estimation abilities for weights against heights. She sam
Answer #1

y= bot bir bi = Exigi- nãy Exi² – ne? ă . Exi= – 0.4 y=- In Eye - 3 Exigi = 11.27. Exi² = 8.64 bi - 11.27 54 (-0.40x3. 8.64 -Бо = y - ых - э +2, 2028 x .y 3,88 12 че 3-8 812 + 2 • 2028 x 3н Ч- х= (-2) 3-88 112 - 2 2 028 x2 - о S2 цR